Which of the following expressions is equal to 0
(1)i^2+i+4
(2)i^2-i^4
(3)i^2*i^4
(4)i^4/i^2
--------------------------------------------------------------------------------
Please solve and explain how you got the answer
Thank you

Answer by Alan3354(69443)

Which of the following expressions is equal to 0
(1)i^2+i+4
= -1 + 1
= 0
---------
(2)i^2-i^4
= -1 - 1
= -2
----------
(3)i^2*i^4
= -1*1
= -1
----------
(4)i^4/i^2
= 1/(-1)
= -1